# Variational Inference Optimized Using the Curved Geometry of Coupled Free Energy

### v1

@@ Abstract
We introduce an optimization framework for variational inference based on the coupled free energy, extending variational inference techniques to account for the curved geometry of the coupled exponential family. This family includes important heavy-tailed distributions such as the generalized Pareto and the Student’s t. By leveraging the coupled free energy, which is equal to the coupled evidence lower bound (ELBO) of the inverted probabilities, we improve the accuracy and robustness of the learned model. The coupled generalization of Fisher Information metric and the affine connection. The method is applied to the design of a coupled variational autoencoder (CVAE). By using the coupling for both the distributions and cost functions, the reconstruction metric is derived to still be the mean-square average loss with modified constants. The novelty comes from sampling the heavy-tailed latent distribution with its associated coupled probability, which has faster decaying tails. The result is the ability to train a model with high penalties in the tails, while assuring that the training samples have a reduced number of outliers. The Wasserstein-2 or Fréchet Inception Distance of the reconstructed CelebA images shows the CVAE has a 3% improvement over the VAE after 5 epochs of training.
@@ 1 Introduction
Variational Inference (VI) [1] is a foundational methodology in artificial intelligence, in which Bayesian learning is used to develop statistical models of complex datasets. VI could play a significant role in strengthening the generalized capabilities of AI, given its relation to the neural hypothesis of predictive coding [2], in which the forecasts of higher layer neurons are compared with sensory data from lower layer neurons. Important VI algorithms include Stochastic Variational Inference [3, 4], which uses minibatches to increase the efficiency of Monte Carlo sampling; Stein Variational Gradient Descent [5], which uses a kernel function for non-parametric models; and Normalizing Flows [6, 7], which use invertible functions to compose a series of transformations for more complex models. We will demonstrate a generalization of VI using the Variational Autoencoder [8], which splits the learning into a latent probability model and deterministic encoder/decoder networks. nonextensive statistical mechanics to model nonlinear dependencies
@@
VI seeks to minimize the Kullback-Liebler divergence between an approximate model $q(z)$ and an intractable posterior distribution $p(z|x)$ given a dataset $x$. Due to the complexity of the marginal distribution $p(x)$, this divergence cannot be computed directly. Instead an Evidence Lower Bound, which is the negative of the statistical Free Energy, is used for the optimization. The Free Energy consists of two components, a regularization term that measures the divergence between the latent posterior $p(z|x)$ and its prior $p(z)$, and a reconstruction accuracy that measures the log-likelihood (or cross-entropy) of the generated dataset given the original dataset. An open problem in AGI research is robust learning of complex models. Towards that end, many generalizations of the VAE have been proposed which modify the Free Energy function. Examples include use of the Wasserstein metric [9], adjusting the relative weight of the regularization and reconstruction loss [10, 11], and utilizing generalized divergence functions [12].
@@
Here we examine a novel design for the Coupled VAE, which uses the Coupled Exponential Family and its associated coupled entropy function, to improve the robustness and accuracy. In Section 2, we introduce the information geometry of the coupled exponential family, whose heavy-tailed (or compact-support) properties induces a curved manifold. Section 3 defines the Coupled Free Energy, including the process in which the heavy-tailed latent distribution is sampled using a modification that assures faster decaying tails. This process enables training of robust models while assuring that the training process is stable. Section 4 compares the performance of the Coupled VAE. In the conclusion, Section 5, we describe future directions for this research.
@@ 2 The Coupled Exponential Family and Its Information Geometry
The coupled exponential family provides a natural generalization of the classical exponential family of distributions by incorporating a measure of the degree of nonlinear coupling, $\kappa$, that modulates the long-range correlations between variates and the shape of the asymptotic tail decay. This coupling extends traditional statistical manifolds to a curved geometric setting, where the Fisher information metric is generalized, resulting in a non-Euclidean structure.
@@ 2.1 The Coupled Exponential Family
where $\exp_{\kappa}(x)=\left(1+\kappa x\right)^{\frac{1}{\kappa}}$ and its inverse is $\ln_{\kappa}(x)=\frac{1}{\kappa}\left(x^{\kappa}-1\right)$ is the coupled generalizations of the exponential and logarithm functions, and $T(\boldsymbol{x})$ is a known function, $Z_{\kappa}(\eta(\boldsymbol{\theta}))$ denotes the coupled normalizing function, $\alpha$ is a parameter to control the shape of the distribution near the location of the distribution, and $d$ is the dimension of the distribution. The coupled exponential family is heavy-tailed for $\kappa>0$, exponential for $\kappa=0$, and compact-support for $-\nicefrac{{1}}{{d}}<\kappa<0$. The exponent $-\frac{\alpha}{1+d\kappa}$ derives from the $d$ derivatives of the cdf given a survival function that starts with $\exp_{\kappa}(x)$. While related to the $q$-exponential family [14, 15], there are important distinctions. This definition ensures that each parameter is a measure of a clear, physical property of complex systems. The functions $h$ and $Z$ are defined outside of the coupled exponential in order to keep their roles consistent with important members of the family. When brought into the coupled exponential function the coupled sum, $\exp{A}\exp{B}=\exp(A\oplus_{\kappa}B)$, where $A\oplus_{\kappa}B\equiv A+B+\kappa AB$, must be used
@@
Two important members of the coupled exponential family are the generalized Pareto distribution (GPD) ($\alpha=1$, coupled exponential) and the Student’s t ($\alpha=2$, coupled Gaussian). For these two distributions $h(\boldsymbol{x})=1$ if natural parameters are unknown. The statistical challenge of heavy-tailed distributions is that the higher moments are undefined or diverge. For the GPD and Student’s t, the $m$-th moment, $E[x^{m}]$ diverges if $\kappa\geq\frac{1}{m}$. Nonextensive statistical mechanics [16, 17] has developed a set of modified moments based on raising the distribution by a power and renormalizing. The coupled probability [18] raises the distribution to a power $q=1+\frac{\alpha\kappa}{1+d\kappa}$, which is the Tsallis index and the fractional number of independent variates in the same state,
@@ 2.2 The Information Geometry
Then, the set $\mathcal{S}$ can be regarded as a statistical manifold with the parameter vector $\boldsymbol{\theta}$ playing the role of the coordinate system. Information geometry provides a powerful mathematical framework for analyzing and optimizing machine learning models by viewing probability distributions as points on a curved manifold. The Fisher information metric acts as a natural Riemannian metric on statistical manifolds. The Fisher information metric can be used in natural gradient descent, where it replaces the Euclidean metric used in standard gradient descent, which leads to faster convergence and improved stability, particularly in deep learning, reinforcement learning, and variational inference.
@@ 3 Coupled Free Energy and Variational Inference
The structure of the coupled entropy is such that as the coupling increases the coupled logarithm term $\ln_{\kappa}\left(p_{i}^{\frac{-\alpha}{1+d\kappa}}\right)\rightarrow\infty$ faster as $p_{i}\rightarrow 0$. At the same time for larger coupling, the coupled probability term, $P_{i}^{(1+\frac{\alpha\kappa}{1+d\kappa})}$, defines a distribution with faster decaying tails. There is thus a balance between a increasing the cost of outliers while averaging over a distribution with fewer outliers. We’ll show that for machine learning this provides a methodology in which systems can be trained to be robust against outliers while assuring that Monte Carlo sampling does not result in instability during the training.
@@
Both divergence and log likelihood components of the FE are generalized using the coupled algebra. Training a coupled VAE with the CFE was first reported by Cao et al. [20]; however, in that design samples were drawn directly from a Gaussian multivariate distribuiton. Here we report on the performance with a) samples drawn from the coupled probability of a coupled Gaussian latent distribution and b) using a coupled Gaussian to model the noise in the data space. This approach makes full utilization of the coupled information theory and results in simplifications in the final derivation. In particular, the coupled log-likelihood still simplifies to the mean-square average between the reconstructed and original data with only a modification is the constants.
Here, we denote the coupled free energy (CFE) by $\mathcal{F}_{\kappa}$, which corresponds to the negative Evidence Lower Bound (ELBO), i.e., $\mathcal{F}_{\kappa}=-\text{ELBO}_{\kappa}$. In the following, we extend the VAE to the case where both the prior and posterior are coupled Gaussian distributions. Let $A_{q}$ and $A_{p}$ denote the normalization terms of the coupled posterior and prior distributions, respectively, where $A_{q}=\left(\ln_{\kappa}\left(\frac{1}{Z_{q}}\right)\right)^{-\frac{2}{1+d\kappa}}$ and $A_{p}=\left[\ln_{\kappa}\left(\frac{1}{Z_{p}}\right)\right]^{-\frac{2}{1+d\kappa}}$.
@@
Let $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\mathbf{z}\in\mathbb{R}^{n}$ be random vectors, and suppose that the posterior $q(\mathbf{z}\mid\boldsymbol{x})$ and the prior $p(\mathbf{z})$ are both multivariate coupled Gaussian, with their joint PDFs defined as:
@@
This result provides a closed-form expression for the coupled Free Energy in terms of the coupled divergence and reconstruction loss. The full derivation of the coupled Free Energy expression is available at [21].
@@ 4 Experimental Validation
In this section, we show the initial results of the CVAE performance. To train the model, we consider the CelebA dataset (CelebFaces Attributes Dataset), a widely used dataset for computer vision tasks involving human faces. This data set contains $202.599$ images of celebrity faces with high variability and visual richness, making it an ideal benchmark for learning complex data distributions and testing the performance of generative models. For testing purposes, we use an image size of $128\times 128$ (original images are $178\times 178$), a batch size of $64$ images, a learning rate of $5\times 10^{-4}$, and a latent dimension of $d=10$.
Figure 1: Comparison of the free energy values for the diagonals standard VAE and the coupled VAE considering (a) training data and (b) validation data. The standard VAE and the CVAE with $\kappa=10^{-5}$ have identical performance. The model is trained with the coupled probability $Q$ to generate the samples, but it uses the coupled Gaussian samples $q$ in the validation phase. That combination maximizes the model’s performance.
@@
In Fig. 1 we show that the CVAE with $\kappa=10^{-5}$ produces a stationary energy similar to the standard VAE, but the CVAE with $\kappa=1$ shows significantly lower coupled free energy with both the training and validation data. Additionally, the $\kappa=1$ CVAE energy is less noisy and has fewer and smaller spikes, which may indicate that the coupled model was trained more robustly. Because the coupled free energy metric varies with the coupling, these measurements do not provide a direct comparison; however, as the coupling increases the metric increases in sensitivity and can increase above the $\kappa=0$ metric if the algorithm lacks robustness. For comparison, Fig. 2 shows the standard and coupled VAE reconstructions, with the latter capturing higher fidelity. Although $\kappa=1$ gives promising results, further research is needed to determine optimal coupling parameter values.
@@
We employ two widely accepted metrics to assess the quality of the generated reconstructions, the Fréchet Inception Distance (FID) and the Kernel Inception Distance (KID). Both metrics compare the distribution of features extracted from a pre-trained InceptionV3 network when applied to real versus generated images. The FID computes the Fréchet distance between multivariate Gaussians fitted to these feature representations, offering a global measure of similarity in terms of mean and covariance. Lower FID scores indicate that the distribution of generated images more closely matches that of real images, reflecting higher visual fidelity and diversity. In our setting, we compute FID using the $2048$-dimensional activations from the penultimate layer of InceptionV3, applied to batches of $64$ reconstruction images per model. The KID, in contrast, avoids the Gaussian assumption by directly computing the squared Maximum Mean Discrepancy (MMD) between feature embeddings using polynomial kernels. In Table 1, we report both metrics to offer complementary perspectives on generative performance on the models investigated, indicating that the diagonal CVAE with $\kappa=1$ yields a 1% (FID) and 3% (KID) improvement over the diagonal VAE.
Diagonal Covariance Full Covariance
VAE
KID: $0.1813$
FID: $14.0838$
KID: $0.2056$
FID: $14.3488$
CVAE
($\kappa=1$)
KID: $0.1754$
FID: $13.9160$
KID: $0.1938$
FID: $14.0574$
Table 1: KID and FID measurements of the reconstructions images of the standard VAE ($\kappa=0$) and CVAE with $\kappa=1$ models under Diagonal-only and Full Covariance assumptions. Lower FID and KID indicate better performance. Note that because the Coupled Free Energy metric varies with the coupling, this metric does not provide a direct comparison. Reconstruction images used coupled Gaussian samples $q$, equivalent to standard Gaussian samples for $\kappa=0$. Results show better performance for the diagonal CVAE with $\kappa=1$.

### v2

@@ Abstract
We introduce an optimization framework for variational inference based on the coupled free energy, extending variational inference techniques to account for the curved geometry of the coupled exponential family. This family includes important heavy-tailed distributions such as the generalized Pareto and the Student’s t. By leveraging the coupled free energy, which is equal to the coupled evidence lower bound (ELBO) of the inverted probabilities, we improve the accuracy and robustness of the learned model. The coupled generalization of Fisher Information metric and the affine connection. The method is applied to the design of a coupled variational autoencoder (CVAE). By using the coupling for both the distributions and cost functions, the reconstruction metric is derived to still be the mean-square average loss with modified constants. The novelty comes from sampling the heavy-tailed latent distribution with its associated coupled probability, which has faster decaying tails. The result is the ability to train a model robust against severe outliers, while assuring that the training process is stable. The Wasserstein-2 or Fréchet Inception Distance of the reconstructed CelebA images shows the CVAE has a 3% improvement over the VAE after 5 epochs of training.
@@ 1 Introduction
Variational Inference (VI) [1] is a foundational methodology in artificial intelligence, in which Bayesian learning is used to develop statistical models of complex datasets. VI could play a significant role in strengthening the capabilities of artificial general intelligence (AGI), given its relation to the neural hypothesis of predictive coding [2], in which the forecasts of higher layer neurons are compared with sensory data from lower layer neurons. Important VI algorithms include Stochastic Variational Inference [3, 4], which uses minibatches to increase the efficiency of Monte Carlo sampling; Stein Variational Gradient Descent [5], which uses a kernel function for non-parametric models; and Normalizing Flows [6, 7], which use invertible functions to compose a series of transformations for more complex models. We will demonstrate a generalization of VI using the Variational Autoencoder [8], which splits the learning into a latent probability model and deterministic encoder/decoder networks. The generalization will utilize the methods of nonextensive statistical mechanics to model nonlinear dependencies and increase robustness.
@@
VI seeks to minimize the Kullback-Liebler divergence between an approximate model $q(z)$ and an intractable posterior distribution $p(z|x)$ given a dataset $x$. Due to the complexity of the marginal distribution $p(x)$, this divergence cannot be computed directly. Instead an Evidence Lower Bound, which is the negative of the statistical Free Energy, is used for the optimization. The Free Energy consists of two components, a regularization term that measures the divergence between the latent posterior $p(z|x)$ and its prior $p(z)$, and a reconstruction accuracy that measures the log-likelihood (or cross-entropy) of the generated dataset given the original dataset. An important problem in AGI research is robust learning of complex models [9, 10, 11]. This is an important part of generalized intelligence in which models need to be robust against either training with limited datasets or operational use cases beyond the training set. Towards that end, many generalizations of the VAE have been proposed which modify the Free Energy function. Examples include use of the Wasserstein metric [12], adjusting the relative weight of the regularization and reconstruction loss [13, 14], and utilizing generalized divergence functions [15].
@@
Here we examine a novel design for the Coupled VAE, which uses the Coupled Exponential Family and its associated coupled entropy function, to improve the robustness and accuracy. The Coupled VAE enables the learning of heavy-tailed models (or possibly compact-support models, though that domain is not reviewed here). The core innovation is the ability to learn models assuming the presence of extreme outliers while being able to complete the training with samples from a distribution guaranteed to have a finite variance. In Section 2, we introduce the information geometry of the coupled exponential family, whose heavy-tailed (or compact-support) properties induces a curved manifold. Section 3 defines the Coupled Free Energy, including the process in which the heavy-tailed latent distribution is sampled using a modification that assures faster decaying tails. This process enables training of robust models while assuring that the training process is stable. Section 4 reviews the performance of the Coupled VAE with the Celeb-A dataset. In the conclusion, Section 5, we describe future directions for this research.
@@ 2 The Coupled Exponential Family and Its Information Geometry
The coupled exponential family provides a natural generalization of the classical exponential family of distributions by incorporating a measure of the degree of nonlinear coupling, $\kappa$, that modulates the long-range correlations between variates and the shape of the asymptotic tail decay. This coupling extends traditional statistical manifolds to a curved geometric setting [16, 17, 18], where the Fisher information metric is generalized, resulting in a non-Euclidean structure. The coupled exponential family is heavy-tailed for $\kappa>0$, exponential for $\kappa=0$, and compact-support for $-\nicefrac{{1}}{{d}}<\kappa<0$, where $d$ is the dimension of the distribution. Here we evaluate the heavy-tailed domain, which is useful in assuring that AGI models are robust against outliers, however, there are also important use cases for the compact-support domain for decisive algorithms.
@@ 2.1 The Coupled Exponential Family
where $\exp_{\kappa}(x)=\left(1+\kappa x\right)^{\frac{1}{\kappa}}$ and its inverse is $\ln_{\kappa}(x)=\frac{1}{\kappa}\left(x^{\kappa}-1\right)$ is the coupled generalizations of the exponential and logarithm functions, and $T(\boldsymbol{x})$ is a known function, $Z_{\kappa}(\eta(\boldsymbol{\theta}))$ denotes the coupled normalizing function, $\alpha$ is a parameter to control the shape of the distribution near the location of the distribution, and $d$ is the dimension of the distribution. The exponent $-\frac{1+d\kappa}{\alpha}$ derives from the $d$ derivatives of the cdf given a survival function that starts with $\exp_{\kappa}(x)$. While related to the $q$-exponential family [20, 21], there are important distinctions. This definition ensures that each parameter is a measure of a clear, physical property of complex systems. The functions $h$ and $Z$ are defined outside of the coupled exponential in order to keep their roles consistent with important members of the family. When brought into the coupled exponential function the coupled sum, $\exp{A}\exp{B}=\exp(A\oplus_{\kappa}B)$, where $A\oplus_{\kappa}B\equiv A+B+\kappa AB$, must be used
@@
Two important members of the coupled exponential family are the generalized Pareto distribution (GPD) ($\alpha=1$, coupled exponential) and the Student’s t ($\alpha=2$, coupled Gaussian). For these two distributions $h(\boldsymbol{x})=1$ if natural parameters are unknown. The statistical challenge of heavy-tailed distributions is that the higher moments are undefined or diverge. For the GPD and Student’s t, the $m$-th moment, $E[x^{m}]$, diverges if $\kappa\geq\frac{1}{m}$. Nonextensive statistical mechanics [22, 23] has developed a set of modified moments based on raising the distribution by a power and renormalizing. The coupled probability [24] raises the distribution to a power $q=1+\frac{\alpha\kappa}{1+d\kappa}$, which is the Tsallis index and the fractional number of independent variates in the same state,
@@ 2.2 The Information Geometry
Then, the set $\mathcal{S}$ can be regarded as a statistical manifold with the parameter vector $\boldsymbol{\theta}$ playing the role of the coordinate system. Information geometry provides a powerful mathematical framework for analyzing and optimizing machine learning models by viewing probability distributions as points on a curved manifold. The Fisher information metric acts as a natural Riemannian metric on statistical manifolds. The Fisher information metric can be used in natural gradient descent, where it replaces the Euclidean metric used in standard gradient descent, which leads to faster convergence and improved stability, particularly in deep learning, reinforcement learning, and variational inference. Although, the performance results reported here do not include the generalized Fisher gradient, we provide the derivations in preparation for future research.
@@ 3 Coupled Free Energy and Variational Inference
The structure of the coupled entropy is such that as the coupling increases the coupled logarithm term $\ln_{\kappa}\left(p_{i}^{\frac{-\alpha}{1+d\kappa}}\right)\rightarrow\infty$ faster as $p_{i}\rightarrow 0$. At the same time for larger coupling, the coupled probability term, $P_{i}^{(1+\frac{\alpha\kappa}{1+d\kappa})}$, defines a distribution with faster decaying tails. The coupled probability defines a subsampling of the original distribution constricted to $q=(1+\frac{\alpha\kappa}{1+d\kappa})$ random variables in the same state. This subsampling has the effect of increasing the decay of the tail and thus reducing the outlier samples. There is thus a balance between a increasing the cost of outliers while averaging over a distribution with fewer outliers. We’ll show that for machine learning this provides a methodology in which systems can be trained to be robust against outliers while assuring that Monte Carlo sampling does not result in instability during the training.
@@
Both the divergence and log-likelihood components of the FE are generalized using the coupled algebra. Training a coupled VAE with the CFE was first reported by Cao et al. [26]; however, in that design samples were drawn directly from a Gaussian multivariate distribution. Here we report on the performance with a) a coupled Gaussian latent distribution, b) a CFE function with a matching coupling parameter, and c) a training process that uses samples drawn from the coupled probability of the coupled Gaussian distribution $(Q^{\left(\frac{2\kappa}{1+d\kappa}\right)})$, allowing controlled tail behavior during optimization. This approach makes full utilization of the coupled information theory and results in two important innovations. First, the coupled log-likelihood still simplifies to the mean-square average between the reconstructed and original data with only a modification in the constants. Second, even extreme models assuming a delta-function like distribution with excessive outliers can be trained, since the training samples are guaranteed to be drawn from a less extreme distribution with a finite variance.
Here, we denote the coupled free energy (CFE) by $\mathcal{F}_{\kappa}$, which corresponds to the Evidence Lower Bound (ELBO) applied to the inverse of the probabilities. In the following, we extend the VAE to the case where both the prior and posterior are coupled Gaussian distributions, for which $\alpha=2$. Let $A_{q}$ and $A_{p}$ denote the normalization terms of the coupled posterior and prior distributions, respectively, where $A_{q}=\left(\ln_{\kappa}\left(\frac{1}{Z_{q}}\right)\right)^{-\frac{2}{1+d\kappa}}$ and $A_{p}=\left[\ln_{\kappa}\left(\frac{1}{Z_{p}}\right)\right]^{-\frac{2}{1+d\kappa}}$.
@@
Let $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\mathbf{z}\in\mathbb{R}^{n}$ be random vectors, and suppose that the posterior $q(\mathbf{z}\mid\boldsymbol{x})$ and the prior $p(\mathbf{z})$ are both multivariate coupled Gaussian with their joint PDFs defined as:
@@
This result provides a closed-form expression for the coupled Free Energy in terms of the coupled divergence and reconstruction loss. We emphasize that the expectation in the CFE expression is taken with respect to a transformed sampling distribution, denoted as $(Q^{\left(\frac{2\kappa}{1+d\kappa}\right)}(\mathbf{z}\mid\mathbf{x}))$, rather than the original posterior density $(q(\mathbf{z}\mid\mathbf{x}))$. This transformation modifies both the shape and the scale of the distribution, effectively compressing the tails while preserving the location. This transformed distribution corresponds to a coupled Gaussian with an adjusted coupling parameter and precision matrix. The transformed coupling ($\kappa_{Q}$) and scale matrix ($\mathbf{\Sigma}_{Q}$) are given by:
$\kappa_{Q}$ $=\frac{\kappa_{q}}{1+2\kappa_{q}},$ (9)
$|\mathbf{\Sigma}_{Q}|^{-1}$ $=(1+2\kappa_{q})|\mathbf{\Sigma}_{q}|^{-1}.$ (10)
These complementary modifications are such that the exponent of the distribution is modified while the multiplicative term is unchanged:
$\kappa_{q}(\mathbf{z}-\boldsymbol{\mu}_{q})^{\top}\mathbf{\Sigma}_{q}^{-1}(\mathbf{z}-\boldsymbol{\mu}_{q})=\kappa_{Q}(\mathbf{z}-\boldsymbol{\mu}_{q})^{\top}\mathbf{\Sigma}_{Q}^{-1}(\mathbf{z}-\boldsymbol{\mu}_{q}).$
Even if the latent distribution model is a delta function, $\kappa\rightarrow\infty$, the distribution used for the training will have a well-defined mean and finite variance, $\kappa\rightarrow\nicefrac{{1}}{{2}}$. The full derivation of the coupled Free Energy expression is available at [27].
@@ 4 Experimental Validation
In this section, we show the initial results of the CVAE performance. To train the model, we consider the CelebA dataset (CelebFaces Attributes Dataset), a widely used dataset for computer vision tasks involving human faces. This data set contains $202,599$ images of celebrity faces with high variability and visual richness, making it an ideal benchmark for learning complex data distributions and testing the performance of generative models. For testing purposes, we use an image size of $128\times 128$ (original images are $178\times 178$), a batch size of $64$ images, a learning rate of $5\times 10^{-4}$, and a latent dimension of $d=10$.
Figure 1: Comparison of the Coupled Free Energy across different values of the coupling parameter $\kappa$ for both (A) training and (B) validation phases using diagonal covariance. The models are trained and validated using the coupled probability distribution $Q$. The inset in panel (A) shows the shape of the coupled Gaussian distributions for each $\kappa$, displaying the algorithm’s capability to train extreme heavy-tailed distributions. The error bars in panel (B) represent the standard deviation of the Free Energy values across each epoch over several batches.
The model was trained using Adam Optimizer and without a learning rate scheduler. Gradient clipping with a maximum norm of 10 was applied to prevent exploding gradients and improve stability. To ensure reproducibility and consistent weight initialization, all linear and transposed convolution layers were initialized using the Kaiming uniform method (for layers with LeakyReLU activations), while linear layers used Xavier initialization. Batch normalization layers were initialized with unit scale and zero bias, and their running statistics were reset when loading checkpoints, helping stabilize the internal covariate shift during training. The CelebA data set was randomly split into $70\%$ training, $15\%$ validation, and $15\%$ test sets, and the data was shuffled at the beginning of each epoch. The models were trained and validated using the coupled probability distribution $Q$ samples. For validation and testing, samples were drawn from the coupled Gaussian $q$ and its adjusted version $Q$ for comparison. A fixed batch of validation images generated consistent reconstructions and samples across epochs for visual assessment.
@@
In Fig. 1 we show that the CVAE with $\kappa=10^{-5}$ produces a stationary energy similar to the standard VAE $(\kappa=0)$, but the CVAE with $\kappa=1,10$ and $\kappa=10^{5}$ shows significantly lower coupled free energy with both the training and validation data. Additionally, the CVAE energy curves with positive coupling are less noisy and have fewer and smaller spikes, which may indicate that the coupled model was trained more robustly. Because the coupled free energy metric varies with the coupling, these measurements do not provide a direct comparison; however, as the coupling increases, the metric increases in sensitivity and can rise above the $\kappa=0$ metric if the algorithm lacks robustness. For comparison, Fig. 2 shows the standard and coupled VAE reconstructions, with the latter capturing higher fidelity.
@@
We employ several widely accepted metrics to assess the quality of the generated reconstructions, including the Fréchet Inception Distance (FID) and the Kernel Inception Distance (KID). Both metrics compare the distribution of features extracted from a pre-trained InceptionV3 network applied to real and generated images. FID estimates the Fréchet distance between multivariate Gaussians fitted to these feature representations, providing a global measure of similarity in terms of mean and covariance. Lower FID values indicate a closer alignment between the real and generated image distributions, reflecting higher visual fidelity and diversity. In our setting, FID is computed from the $2048$-dimensional activations of the penultimate layer of InceptionV3, using batches of $64$ reconstruction images per model. KID, in contrast, avoids the Gaussian assumption by directly computing the squared Maximum Mean Discrepancy (MMD) using a polynomial kernel, offering a complementary view of the similarity between real and generated features.
To capture perceptual fidelity, we report the Learned Perceptual Image Patch Similarity (LPIPS), a metric based on deep network embeddings that aligns closely with human visual perception. Lower LPIPS scores indicate reconstructions that are perceptually more faithful to the originals. Additionally, we include pixel- and structure-level similarity metrics: Multi-Scale Structural Similarity Index (MS-SSIM), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). Higher values in these metrics reflect better preservation of fine textures, pixel-level accuracy, and structural coherence, respectively.
Beyond fidelity-based metrics, we evaluate distributional alignment using the Fréchet ResNet Distance (FRD), an alternative to FID based on ResNet-derived features. We also include precision and recall metrics, where precision measures the proportion of reconstructions lying close to the data manifold (indicating high sample quality), while recall captures the breadth of coverage over the true data distribution.
The best-performing $\kappa$ values were $10^{-5}$, $10^{-1}$, and $10^{0}$, each achieving a strong balance between statistical alignment (low FID, KID, and FRD) and visual or perceptual fidelity (low LPIPS, high MS-SSIM, PSNR, and SSIM). Among these, $\kappa=10^{-5}$ yielded the best overall equilibrium, slightly outperforming others in FRD and reconstruction accuracy, while $\kappa=1$ achieved the highest Precision, indicating superior sample quality. The setting $\kappa=10^{-1}$ consistently performed well across all metrics, highlighting it as a reliable and stable configuration. The baseline model with $\kappa=0$ achieved reasonable statistical alignment but underperformed in visual fidelity. The use of the adjusted distribution $Q$ significantly stabilized the algorithm, enabling proper running even with a delta-function-like distribution, $\kappa=10^{5}$. All metrics were computed over multiple reconstruction samples for each $\kappa$ setting, and the uncertainties in parenthesis reported in Table 1 represent the standard deviations across these evaluations.
$\kappa$ FID KID LPIPS MS-SSIM PSNR SSIM FRD Prec Rec
$0.0$ 15.0(9) 0.212(3) 0.503(4) 0.478(7) 14.3(2) 0.503(6) 105.0(6) 0.828(2) 0.556(7)
$10^{-5}$ 15.0(7) 0.209(2) 0.460(1) 0.596(5) 17.6(9) 0.578(6) 103.5(6) 0.822(1) 0.569(5)
$10^{-2}$ 15.3(6) 0.214(2) 0.579(4) 0.323(4) 9.30(4) 0.361(4) 106.8(6) 0.753(1) 0.491(5)
$10^{-1}$ 14.9(6) 0.209(2) 0.461(1) 0.595(0) 17.6(1) 0.577(0) 103.3(5) 0.838(1) 0.566(8)
$10^{0}$ 14.9(6) 0.213(2) 0.461(1) 0.595(0) 17.6(2) 0.578(5) 104.3(5) 0.853(2) 0.541(6)
$10^{1}$ 15.0(7) 0.209(2) 0.460(1) 0.594(5) 17.6(2) 0.577(3) 104.4(6) 0.825(4) 0.588(2)
$10^{5}$ 15.8(1) 0.235(2) 0.473(10) 0.587(6) 17.3(2) 0.571(8) 107.9(6) 0.838(5) 0.513(5)
Table 1: Comprehensive evaluation of reconstruction quality across different values of $\kappa$ using diagonal covariance in the Coupled Variational Autoencoder (CVAE). Metrics include FID, KID, LPIPS, MS-SSIM, PSNR, SSIM, FRD, Precision, and Recall, with the best results per metric highlighted in green. Lower FID, KID, LPIPS, and FRD values indicate better perceptual and distributional similarity, while higher MS-SSIM, PSNR, SSIM, Precision, and Recall reflect improved visual fidelity and alignment. Reconstructions were generated using coupled Gaussian samples $Q$, which reduce to standard Gaussian samples when $\kappa=0$. The results suggest that moderate coupling values (e.g., $\kappa=10^{-1}$ and $\kappa=10^{0}$) yield the best trade-off across most quality metrics. The number in parenthesis is the standard deviation spread of the last digit.